OpenStudy (anonymous):

fan-ing and medal if you answer? :3

OpenStudy (anonymous):

The average annual salary of the employees of a company in the year 2005 was $80 thousand. It increased by the same factor each year and in 2006, the average annual salary was $88 thousand. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)? f(x) = 88(0.88)x f(x) = 88(1.1)x f(x) = 80(0.88)x f(x) = 80(1.1)x

OpenStudy (kc_kennylau):

What's the ratio(r)? What's the first term(a)? f(x) = ar^x

OpenStudy (anonymous):

f(x) = 80(1.1)x

OpenStudy (anonymous):

f(1)=80(1.1)1

OpenStudy (anonymous):

Thx :)